# Quantum Machine Learning: from NISQ to Fault Tolerance

### v1

@@ IV.1 Shadow tomography
Shadow tomography was a proposal first came up with by Scott Aaronson couple of years ago Aaronson (2018). In this final section, we will review some key points of it and also a little about its recent developments.
@@
The complexity increases significantly when dealing with $n$ qubits. For a pure state of $n$ qubits, obtaining an approximate description requires dealing with $2^{n}$ complex numbers. This exponential growth poses a challenge, as it would necessitate exponential classical bits to represent $n$ qubits. Even for a relatively modest system of 1000 qubits, it would require $2^{1000}$ complex numbers for an approximate description. To put this into perspective, this exceeds the total number of atoms in the observable universe, estimated to be around $10^{80}$. This realization, first acknowledged in the 1980s, played a crucial role in the conceptualization of quantum computing by visionaries like Richard Feynman Feynman et al. (1996) and others. But again measurement will yield at most $n$ bits.
@@
The quantum state tomography challenge involves a machine capable of generating $n$-qubit quantum states and producing identical copies of a specified state. Focusing on a mixed quantum state represented by a $D\times D$ density matrix, denoted as $\rho$, the goal is to gain insights into this density matrix. The key question is to determine the minimum number of copies of the state $\rho$ necessary to effectively tackle this challenge. The answer to this question is approximately $\mathcal{O}(D^{2})$ copies of $\rho$, which was established in 2016 O’Donnell and Wright (2015). This implies that the number of required copies scales polynomially with $D$, the dimension of the Hilbert space, which is still exponentially large. For instance, learning about 100 qubits would necessitate around $2^{200}$ copies of the state, an impractical and astronomically large number. Consequently, the feasibility of quantum state tomography becomes increasingly limited as the number of qubits grows, and one of the world records stands at about 10 qubits Song et al. (2017), involving millions of measurement settings. This underscores the significant challenges and limitations in quantum state tomography for larger quantum systems.
@@
The concept of shadow tomography, introduced by Scott Aaronson Aaronson (2018, 2020), is motivated by the following scenario. We are provided with an unknown $D$-dimensional mixed state $\rho$ and a set of observables, $\mathcal{O}_{1},\ldots,\mathcal{O}_{M}$. Assuming these observables can only take two values, such as yes or no, and considering that these measurements fully characterize the state, we aim to recover the “shadow” cast by $\rho$ on the measurements $\mathcal{O}_{1},\ldots,\mathcal{O}_{M}$. The term “shadow tomography” was suggested by Steve Flammia, emphasizing our goal of retrieving the “shadow” rather than the entire density matrix of $\rho$. The objective is to determine the minimum number of states $\rho$ needed to estimate the expectation value $\operatorname{Pr}[\mathcal{O}_{i}]=\operatorname{Tr}(\mathcal{O}_{i}\rho)$ for all these observables.
@@
Our intuition suggests that approximately $\mathcal{O}(D^{2})$ copies would be adequate. We can bypass the entire shadow tomography setup and simply obtain a well-approximated version of the state $\rho$. Once we possess the state, accurately determining any expectation values becomes feasible. Furthermore, it seems intuitive that having roughly $\mathcal{O}(M)$ copies of the state should enable straightforward shadow tomography. By conducting measurements on all $M$ observables, we can derive the expectation values. Although the state is lost after each measurement, the availability of numerous copies remains. The latter will be practical if $M$ is small, which means that we only care a small number of measurements or observables. But in many cases, both $D$ and $M$ are enormous. So the crucial question revolves around whether there exist a systematic way of accomplishing shadow tomography can be accomplished with a time complexity of $\operatorname{poly}(\log{D},\log{M})~{}.$
@@
In 2018, it was demonstrated in Aaronson (2018) that a protocol exists capable of achieving shadow tomography with a time complexity of $\operatorname{poly}(\log{D},\log{M})$.
Theorem IV.1.
@@
Shadow tomography is solvable using only $\tilde{\mathcal{O}}\left(\log^{4}{M}\cdot\log{D}/\varepsilon^{4}\right)$ copies of the state $\rho~{},$ where the $\tilde{\mathcal{O}}$ hides a higher order factor. The procedure is fully explicit.
@@
The implications of Theorem IV.1 include the following:
@@
To comprehend the workings of shadow tomography, a key concept is “gentle measurement” Aaronson and Rothblum (2019). If a measurement of a mixed state $\rho$ yields a certain outcome with a probability greater than or equal to $1-\varepsilon$, then after the measurement, there exists a state $\rho^{\prime}$ that is $\sqrt{\varepsilon}$-close to $\rho$ in trace distance. In simpler terms, if a state is close to an eigenstate of some observable, it is highly likely that the state will essentially collapse into that eigenstate, which isn’t much different from the original state before the measurement occurred. An even stronger statement holds: applying $M$ such measurements in succession, each accepting with a probability greater than or equal to $1-2M\sqrt{\varepsilon}$, ensures that the error grows linearly. Thus, if a sequence of measurements is applied, and each one does not significantly damage $\rho$, the cumulative effect of the entire session (assuming not too many measurements, i.e., small $M$) will also not significantly damage $\rho$.
@@
One obvious question might be why gentle measurement does not solve our problem of shadow tomography immediately. Gentle measurement only provides a solution under a special case known as a **promise gap**. The gentle measurement will give us an answer under the following circumstance: given measurements $\mathcal{O}_{1},\ldots,\mathcal{O}_{M}$ and real numbers $C_{1},\ldots,C_{M}$, and being promised that $\operatorname{Pr}[\mathcal{O}_{i}\text{ accepts }\rho]$ is either at least $C_{i}$ or at most $C_{i}-\varepsilon$ for each $i$, we can decide for each $i$ using only $\mathcal{O}\left(\log{M}/\varepsilon^{2}\right)$ copies of $\rho$. This provides a good zone for gentle measurements to figure out which side of the two outcome measurements our state is close to, allowing gentle measurement to work. The problem arises when we are in the intermediate zone: $C_{i}-\varepsilon\geq\operatorname{Pr}[\mathcal{O}_{i}\text{ accepts }\rho]\geq C_{i}$. In this case, we will not be able to use this trick and accomplish shadow tomography. In other words, when there’s no promise, we can never rule out that we’re on the knife-edge between acceptance and rejection, making measuring $\rho$ dangerous, since measurement there will not be gentle and destroy our state. And because our objective limits our copies of $\rho$ to only $\log{M}$, we can’t afford to destroy that many copies of the state.
The whole point of shadow tomography is to address this problem Aaronson (2018, 2020). The actual proof of Theorem IV.1 involves a combination of several tricks. The first one was intended to solve a different yet similar problem. Imagine Alice, who knows the state $\rho$, needs to send its information or description to Bob with respect to its behavior on measurements of observables $\mathcal{O}_{1},\ldots,\mathcal{O}_{M}$, which we call $E_{1},\ldots,E_{M}$. In contrast to shadow tomography, Alice does know the state at the beginning, but she has to send the information about the measurements in much fewer bits.
@@
Another protocol to be combined with the previous one is called the Quantum OR Bound Harrow et al. (2017).
Theorem IV.2.
@@
Let $\rho$ be an unknown mixed state, and let $E_{1}~{},\ldots~{},E_{M}$ be known 2-outcome measurements. Suppose we are promised that either
@@
    there exists an $i$ such that $\operatorname{Pr}[E_{i}\text{ accepts }\rho]\geq C~{};$ or else
@@
    $\operatorname{Pr}[E_{i}\text{ accepts }\rho]\leq C-\varepsilon~{},$  for all $i\in[M]~{}.$
@@
Then we can decide which, with high probability, given only $\mathcal{O}\left(\log{M}/\varepsilon^{2}\right)$ copies of $\rho~{}.$
@@
So what’s new about shadow tomography is the combination of two protocols, postselection learning and the Quantum OR Bound. Scott Aaronson Aaronson (2018) found that you can use the Quantum OR Bound protocol as a subroutine to repeatedly search for informative measurements. Again, you start with the maximally mixed state and use the quantum OR bound repeatedly to search for measurements that can best improve your guess, round by round, to shape it into the one that has a similar outcome in terms of measurements with the target state $\rho$. Putting everything together, one will end up with Theorem IV.1.
@@
The result in Theorem IV.1 is not the upper or lower bound; it simply shows that shadow tomography can be done in $\log{M}$ steps. Nowadays, the best lower bound people can prove is $\mathcal{O}(\log{M}/\varepsilon^{2})$, which holds even if we are trying to learn a classical distribution Aaronson and Rothblum (2019). So it is still unclear whether we need quantum Hilbert dimension $D$ in Theorem IV.1. In Aaronson (2020), the authors were able to improve upon the $\log{M}$ dependence using a brand-new connection between gentle measurement and differential privacy. The sample complexity they have achieved is $k=\mathcal{O}\left(\log^{2}{M}\cdot\log^{2}{D}/\varepsilon^{8}\right)~{}.$
@@
Inspired by shadow tomography, Huang et al. (2021a) introduced the classical shadow protocol, as discussed in section II.1.3. Thus, quantum machine learning is a rapidly developing field that encompasses various topics in physics and computer science. Although sub-fields may appear distinct, they all fall under the umbrella of quantum machine learning, influencing and potentially finding applications in one another. As a result, as Scott Aaronson states at the end of his paper Aaronson (2015), despite decades of research in quantum computing, researchers still marvel at the fact that the laws of quantum physics enable us to solve classical problems significantly faster than today’s computers seem capable of. Therefore, it should perhaps not surprise us that, in machine learning and elsewhere, unlocking the full potential of quantum speedups requires significant effort.
IV.2 Classical shadow formalism and random measurements
@@
Drawing inspiration from the principles of shadow tomography, the classical shadow formalism has emerged as an intriguing topic with the potential to serve as an efficient quantum-classical information converter. A series of works, such as Huang et al. (2020b, 2021a); Elben et al. (2022), have contributed to establishing a rigorous framework for reasoning about the randomized measurement paradigm, a key component of classical shadow. These works not only delve into the conceptual aspects but also provide methods to derive error bounds for quantum information extraction based on randomized measurements, leveraging probability theory.
@@
The challenges associated with the quantum-classical information interface, particularly in the context of NISQ algorithms (discussed in sections II.1 and II.2), as well as in the FTQC era for quantum machine learning algorithms (explored in section III.3.1), highlight the necessity of efficient readout mechanisms. Researchers are actively addressing the need for reliable ways to input and extract information from quantum systems to classical ones, as the success of quantum machine learning programs relies on overcoming these challenges. If the process of loading classical data into a quantum computer or transforming quantum data into classical form takes an significant amount of time, the promised speed-ups of quantum algorithms become impractical. Therefore, not only is it crucial to find methods for transforming classical data into quantum form and vice versa, but the efficiency of these processes, including achieving significant speed-up, is paramount, given the exponential growth of Hilbert space degrees of freedom.
@@
The most direct and comprehensive solution to the input problem is Quantum Random Access Memory (QRAM), as detailed in section III.4. Additionally, alternative techniques like pruning, discussed in section III.3.1, offer potential solutions to circumvent the pressing requirement for QRAM. Classical shadow, see also section II.1.3, on the other hand, is focused on addressing the output problem. In this scenario, we possess a quantum state that encapsulates information trained by our Quantum Machine Learning (QML) program, and the goal is to decode this quantum state into classical information for straightforward interpretation.
@@
To present this challenge in a broader context, the task is to efficiently estimate the expectation values of multiple structured observables. Specifically, the aim is to estimate the expectation value of various individual terms in Hamiltonians, where these terms share a well-defined structure, such as being Pauli operators. The proposed strategy for addressing this challenge involves diverting attention from the intricate details of the terms and instead uniformly sampling Pauli measurements. In simpler terms, the goal is to estimate the expectation values of Pauli observables, with the selection of the Pauli basis for each qubit being determined randomly during the measurement process Elben et al. (2022).
@@
This concept can be grasped intuitively through the lens of the infinite monkey theorem Evans et al. (2019). The theorem posits that a monkey randomly hitting keys on a typewriter for an infinite amount of time will almost certainly type any given text, including the complete works of William Shakespeare, and will likely produce every possible finite text an infinite number of times. Let’s apply this idea to predicting an arbitrary low-weight Pauli observable, represented by the purple blocks in Fig. 9. The white blocks signify identity matrices, which are not of primary concern. The crucial aspect is the presence of three specific Pauli operators adjacent to each other in Fig. 9. Now, if we randomly assign Pauli measurements to each qubit, we are free to choose any Pauli measurements for the white blocks on the left side of Fig. 9. However, for the purple blocks, we must use the same Pauli measurements as on the left, or else we won’t gain meaningful information about the statistics. To achieve this, random sampling is employed. Importantly, the performance is not adversely affected, as the probability of obtaining such a low-weight Pauli string is exponentially small, but this exponentiation is in the weight of the Pauli, not in the number of qubits Evans et al. (2019). Additionally, if single-shot measurements are considered, there will be an oversampling factor.
@@
The crucial aspect is that, since everything is done randomly, an effective union bound can be applied to estimate numerous low-weight Pauli observables. The union bound scales logarithmically with the number of terms. Consequently, if one randomly samples each Pauli measurement, the expected total number of measurements required to predict $L$ low-weight observables with accuracy $\epsilon$ each is only logarithmic in $L$, thanks to the union bound Elben et al. (2022).
To formally understand the classical shadow formalism, first we construct a quantum channel which transforms the quantum data into the classical data. We start with an unknown $m$ qubit quantum state $\rho$ and do a certain random unitary rotation: $U~{},$ for example an independent single qubit Clifford rotation. Then we do a computational basis measurement on all of them. Since we did measurement at last, so the output of this channel is actually, honest to god, classical information. This, in Fig. 10, means that $U^{\dagger}|b\rangle\langle b|U$ although written in quantum language, is classical data. Hence, altogether, what this channel does is to transform the quantum state we got from the quantum computer and mapped it into classical information which we will do further computation.
@@
The subsequent step involves averaging the random unitary sampling (through an integral over $U$), depending on the chosen ensemble, and also averaging over the computational basis outcomes (by summing over $b$) as per Born’s rule. If the ensemble is well-behaved, such as Haar random, then the result after the averaging step can be efficiently computed. Moreover, since the result is classical information rather than quantum, it can be easily inverted to obtain the state $\rho$. Consequently, the state $\rho$ is obtained by performing random measurements to acquire classical data, taking average, and then inverting the result to retrieve the state. The lingering question pertains to the convergence speed of this program to the desired expectation value. Drawing inspiration from shadow tomography, the authors of Huang et al. (2020a) demonstrated that the number of required snapshots or single measurements in this program scales logarithmically with the number of terms needed to predict: $N\gtrsim\log{(L)}\max_{j}||\mathcal{O}_{j}||_{\text{shadow}}^{2}/\epsilon^{2}$, where $\mathcal{O}_{j}$ is the observable to be measured, and $\epsilon$ denotes the desired accuracy.
@@
In 2021, the authors of Struchalin et al. (2021) conducted an experimental implementation of classical shadow, demonstrating its superiority over maximum likelihood in the low sampling regime. This framework’s applicability extends to the prediction of observable polynomials, as discussed in Elben et al. (2022). The strategy involves expressing polynomials as linear combinations of tensor products and substituting the tensor products with independent shadows, enhancing efficiency in probing entanglements in many-body systems. Notably, this method exhibits robustness in the presence of noise, leveraging random unitaries as detailed in Chen et al. (2021). Additionally, Huang et al. (2021c) explores techniques for derandomizing the procedure. In classical computer science, there exists a standard approach to transform a probabilistic argument into a deterministic strategy, maintaining comparable effectiveness without reliance on randomness.
IV.2.1 Machine learning with classical shadow

### v2

@@ 4.1 Shadow tomography
Shadow tomography was a proposal first came up with by Scott Aaronson couple of years ago [2]. In this final section, we will review some key points of it and also a little about its recent developments.
@@
The complexity increases significantly when dealing with $n$ qubits. For a pure state of $n$ qubits, obtaining an approximate description requires dealing with $2^{n}$ complex numbers. This exponential growth poses a challenge, as it would necessitate exponential classical bits to represent $n$ qubits. Even for a relatively modest system of 1000 qubits, it would require $2^{1000}$ complex numbers for an approximate description. To put this into perspective, this exceeds the total number of atoms in the observable universe, estimated to be around $10^{80}$. This realization, first acknowledged in the 1980s, played a crucial role in the conceptualization of quantum computing by visionaries like Richard Feynman [45] and others. But again measurement will yield at most $n$ bits.
@@
The quantum state tomography challenge involves a machine capable of generating $n$-qubit quantum states and producing identical copies of a specified state. Focusing on a mixed quantum state represented by a $D\times D$ density matrix, denoted as $\rho$, the goal is to gain insights into this density matrix. The key question is to determine the minimum number of copies of the state $\rho$ necessary to effectively tackle this challenge. The answer to this question is approximately $\mathcal{O}(D^{2})$ copies of $\rho$, which was established in 2016 [104]. This implies that the number of required copies scales polynomially with $D$, the dimension of the Hilbert space, which is still exponentially large. For instance, learning about 100 qubits would necessitate around $2^{200}$ copies of the state, an impractical and astronomically large number. Consequently, the feasibility of quantum state tomography becomes increasingly limited as the number of qubits grows, and one of the world records stands at about 10 qubits [118], involving millions of measurement settings. This underscores the significant challenges and limitations in quantum state tomography for larger quantum systems.
@@
The concept of shadow tomography, introduced by Scott Aaronson [2, 3], is motivated by the following scenario. We are provided with an unknown $D$-dimensional mixed state $\rho$ and a set of observables, $\mathcal{O}_{1},\ldots,\mathcal{O}_{M}$. Assuming these observables can only take two values, such as yes or no, and considering that these measurements fully characterize the state, we aim to recover the “shadow” cast by $\rho$ on the measurements $\mathcal{O}_{1},\ldots,\mathcal{O}_{M}$. The term “shadow tomography” was suggested by Steve Flammia, emphasizing our goal of retrieving the “shadow” rather than the entire density matrix of $\rho$. The objective is to determine the minimum number of states $\rho$ needed to estimate the expectation value $\operatorname{Pr}[\mathcal{O}_{i}]=\operatorname{Tr}(\mathcal{O}_{i}\rho)$ for all these observables.
@@
Our intuition suggests that approximately $\mathcal{O}(D^{2})$ copies would be adequate. We can bypass the entire shadow tomography setup and simply obtain a well-approximated version of the state $\rho$. Once we possess the state, accurately determining any expectation values becomes feasible. Furthermore, it seems intuitive that having roughly $\mathcal{O}(M)$ copies of the state should enable straightforward shadow tomography. By conducting measurements on all $M$ observables, we can derive the expectation values. Although the state is lost after each measurement, the availability of numerous copies remains. The latter will be practical if $M$ is small, which means that we only care a small number of measurements or observables. But in many cases, both $D$ and $M$ are enormous. So the crucial question revolves around whether there exist a systematic way of accomplishing shadow tomography can be accomplished with a time complexity of $\operatorname{poly}(\log{D},\log{M})\leavevmode\nobreak\ .$
@@
In 2018, it was demonstrated in [2] that a protocol exists capable of achieving shadow tomography with a time complexity of $\operatorname{poly}(\log{D},\log{M})$.
Theorem 4.1.
@@
Shadow tomography is solvable using only $\tilde{\mathcal{O}}\left(\log^{4}{M}\cdot\log{D}/\varepsilon^{4}\right)$ copies of the state $\rho\leavevmode\nobreak\ ,$ where the $\tilde{\mathcal{O}}$ hides a higher order factor. The procedure is fully explicit.
@@
The implications of Theorem 4.1 include the following:
@@
To comprehend the workings of shadow tomography, a key concept is “gentle measurement” [4]. If a measurement of a mixed state $\rho$ yields a certain outcome with a probability greater than or equal to $1-\varepsilon$, then after the measurement, there exists a state $\rho^{\prime}$ that is $\sqrt{\varepsilon}$-close to $\rho$ in trace distance. In simpler terms, if a state is close to an eigenstate of some observable, it is highly likely that the state will essentially collapse into that eigenstate, which isn’t much different from the original state before the measurement occurred. An even stronger statement holds: applying $M$ such measurements in succession, each accepting with a probability greater than or equal to $1-2M\sqrt{\varepsilon}$, ensures that the error grows linearly. Thus, if a sequence of measurements is applied, and each one does not significantly damage $\rho$, the cumulative effect of the entire session (assuming not too many measurements, i.e., small $M$) will also not significantly damage $\rho$.
@@
One obvious question might be why gentle measurement does not solve our problem of shadow tomography immediately. Gentle measurement only provides a solution under a special case known as a promise gap. The gentle measurement will give us an answer under the following circumstance: given measurements $\mathcal{O}_{1},\ldots,\mathcal{O}_{M}$ and real numbers $C_{1},\ldots,C_{M}$, and being promised that $\operatorname{Pr}[\mathcal{O}_{i}\text{ accepts }\rho]$ is either at least $C_{i}$ or at most $C_{i}-\varepsilon$ for each $i$, we can decide for each $i$ using only $\mathcal{O}\left(\log{M}/\varepsilon^{2}\right)$ copies of $\rho$. This provides a good zone for gentle measurements to figure out which side of the two outcome measurements our state is close to, allowing gentle measurement to work. The problem arises when we are in the intermediate zone: $C_{i}-\varepsilon\geq\operatorname{Pr}[\mathcal{O}_{i}\text{ accepts }\rho]\geq C_{i}$. In this case, we will not be able to use this trick and accomplish shadow tomography. In other words, when there’s no promise, we can never rule out that we’re on the knife-edge between acceptance and rejection, making measuring $\rho$ dangerous, since measurement there will not be gentle and destroy our state. And because our objective limits our copies of $\rho$ to only $\log{M}$, we can’t afford to destroy that many copies of the state.
The whole point of shadow tomography is to address this problem [2, 3]. The actual proof of Theorem 4.1 involves a combination of several tricks. The first one was intended to solve a different yet similar problem. Imagine Alice, who knows the state $\rho$, needs to send its information or description to Bob with respect to its behavior on measurements of observables $\mathcal{O}_{1},\ldots,\mathcal{O}_{M}$, which we call $E_{1},\ldots,E_{M}$. In contrast to shadow tomography, Alice does know the state at the beginning, but she has to send the information about the measurements in much fewer bits.
@@
Another protocol to be combined with the previous one is called the Quantum OR Bound [59].
Theorem 4.2.
@@
Let $\rho$ be an unknown mixed state, and let $E_{1}\leavevmode\nobreak\ ,\ldots\leavevmode\nobreak\ ,E_{M}$ be known 2-outcome measurements. Suppose we are promised that either
@@
    there exists an $i$ such that $\operatorname{Pr}[E_{i}\text{ accepts }\rho]\geq C\leavevmode\nobreak\ ;$ or else
@@
    $\operatorname{Pr}[E_{i}\text{ accepts }\rho]\leq C-\varepsilon\leavevmode\nobreak\ ,$  for all $i\in[M]\leavevmode\nobreak\ .$
@@
Then we can decide which, with high probability, given only $\mathcal{O}\left(\log{M}/\varepsilon^{2}\right)$ copies of $\rho\leavevmode\nobreak\ .$
@@
So what’s new about shadow tomography is the combination of two protocols, postselection learning and the Quantum OR Bound. Scott Aaronson [2] found that you can use the Quantum OR Bound protocol as a subroutine to repeatedly search for informative measurements. Again, you start with the maximally mixed state and use the quantum OR bound repeatedly to search for measurements that can best improve your guess, round by round, to shape it into the one that has a similar outcome in terms of measurements with the target state $\rho$. Putting everything together, one will end up with Theorem 4.1.
@@
The result in Theorem 4.1 is not the upper or lower bound; it simply shows that shadow tomography can be done in $\log{M}$ steps. Nowadays, the best lower bound people can prove is $\mathcal{O}(\log{M}/\varepsilon^{2})$, which holds even if we are trying to learn a classical distribution [4]. So it is still unclear whether we need quantum Hilbert dimension $D$ in Theorem 4.1. In [3], the authors were able to improve upon the $\log{M}$ dependence using a brand-new connection between gentle measurement and differential privacy. The sample complexity they have achieved is $k=\mathcal{O}\left(\log^{2}{M}\cdot\log^{2}{D}/\varepsilon^{8}\right)\leavevmode\nobreak\ .$
@@
Inspired by shadow tomography, [66] introduced the classical shadow protocol, as discussed in section 2.1.3. Thus, quantum machine learning is a rapidly developing field that encompasses various topics in physics and computer science. Although sub-fields may appear distinct, they all fall under the umbrella of quantum machine learning, influencing and potentially finding applications in one another. As a result, as Scott Aaronson states at the end of his paper [1], despite decades of research in quantum computing, researchers still marvel at the fact that the laws of quantum physics enable us to solve classical problems significantly faster than today’s computers seem capable of. Therefore, it should perhaps not surprise us that, in machine learning and elsewhere, unlocking the full potential of quantum speedups requires significant effort.
4.2 Classical shadow formalism and random measurements
@@
Drawing inspiration from the principles of shadow tomography, the classical shadow formalism has emerged as an intriguing topic with the potential to serve as an efficient quantum-classical information converter. A series of works, such as [63, 66, 41], have contributed to establishing a rigorous framework for reasoning about the randomized measurement paradigm, a key component of classical shadow. These works not only delve into the conceptual aspects but also provide methods to derive error bounds for quantum information extraction based on randomized measurements, leveraging probability theory.
@@
The challenges associated with the quantum-classical information interface, particularly in the context of NISQ algorithms (discussed in sections 2.1 and 2.2), as well as in the FTQC era for quantum machine learning algorithms (explored in section 3.3.1), highlight the necessity of efficient readout mechanisms. Researchers are actively addressing the need for reliable ways to input and extract information from quantum systems to classical ones, as the success of quantum machine learning programs relies on overcoming these challenges. If the process of loading classical data into a quantum computer or transforming quantum data into classical form takes an significant amount of time, the promised speed-ups of quantum algorithms become impractical. Therefore, not only is it crucial to find methods for transforming classical data into quantum form and vice versa, but the efficiency of these processes, including achieving significant speed-up, is paramount, given the exponential growth of Hilbert space degrees of freedom.
@@
The most direct and comprehensive solution to the input problem is Quantum Random Access Memory (QRAM), as detailed in section 3.4. Additionally, alternative techniques like pruning, discussed in section 3.3.1, offer potential solutions to circumvent the pressing requirement for QRAM. Classical shadow, see also section 2.1.3, on the other hand, is focused on addressing the output problem. In this scenario, we possess a quantum state that encapsulates information trained by our Quantum Machine Learning (QML) program, and the goal is to decode this quantum state into classical information for straightforward interpretation.
@@
To present this challenge in a broader context, the task is to efficiently estimate the expectation values of multiple structured observables. Specifically, the aim is to estimate the expectation value of various individual terms in Hamiltonians, where these terms share a well-defined structure, such as being Pauli operators. The proposed strategy for addressing this challenge involves diverting attention from the intricate details of the terms and instead uniformly sampling Pauli measurements. In simpler terms, the goal is to estimate the expectation values of Pauli observables, with the selection of the Pauli basis for each qubit being determined randomly during the measurement process [41].
@@
This concept can be grasped intuitively through the lens of the infinite monkey theorem [43]. The theorem posits that a monkey randomly hitting keys on a typewriter for an infinite amount of time will almost certainly type any given text, including the complete works of William Shakespeare, and will likely produce every possible finite text an infinite number of times. Let’s apply this idea to predicting an arbitrary low-weight Pauli observable, represented by the purple blocks in Fig. 9. The white blocks signify identity matrices, which are not of primary concern. The crucial aspect is the presence of three specific Pauli operators adjacent to each other in Fig. 9. Now, if we randomly assign Pauli measurements to each qubit, we are free to choose any Pauli measurements for the white blocks on the left side of Fig. 9. However, for the purple blocks, we must use the same Pauli measurements as on the left, or else we won’t gain meaningful information about the statistics. To achieve this, random sampling is employed. Importantly, the performance is not adversely affected, as the probability of obtaining such a low-weight Pauli string is exponentially small, but this exponentiation is in the weight of the Pauli, not in the number of qubits [43]. Additionally, if single-shot measurements are considered, there will be an oversampling factor.
@@
The crucial aspect is that, since everything is done randomly, an effective union bound can be applied to estimate numerous low-weight Pauli observables. The union bound scales logarithmically with the number of terms. Consequently, if one randomly samples each Pauli measurement, the expected total number of measurements required to predict $L$ low-weight observables with accuracy $\epsilon$ each is only logarithmic in $L$, thanks to the union bound [41].
To formally understand the classical shadow formalism, first we construct a quantum channel which transforms the quantum data into the classical data. We start with an unknown $m$ qubit quantum state $\rho$ and do a certain random unitary rotation: $U\leavevmode\nobreak\ ,$ for example an independent single qubit Clifford rotation. Then we do a computational basis measurement on all of them. Since we did measurement at last, so the output of this channel is actually, honest to god, classical information. This, in Fig. 10, means that $U^{\dagger}|b\rangle\langle b|U$ although written in quantum language, is classical data. Hence, altogether, what this channel does is to transform the quantum state we got from the quantum computer and mapped it into classical information which we will do further computation.
@@
The subsequent step involves averaging the random unitary sampling (through an integral over $U$), depending on the chosen ensemble, and also averaging over the computational basis outcomes (by summing over $b$) as per Born’s rule. If the ensemble is well-behaved, such as Haar random, then the result after the averaging step can be efficiently computed. Moreover, since the result is classical information rather than quantum, it can be easily inverted to obtain the state $\rho$. Another example may be found in [23], where the authors consider classical shadows with unitary ensembles that are invariant under multiplication by Pauli operators. Consequently, the state $\rho$ is obtained by performing random measurements to acquire classical data, taking average, and then inverting the result to retrieve the state. The lingering question pertains to the convergence speed of this program to the desired expectation value. Drawing inspiration from shadow tomography, the authors of [62] demonstrated that the number of required snapshots or single measurements in this program scales logarithmically with the number of terms needed to predict: $N\gtrsim\log{(L)}\max_{j}||\mathcal{O}_{j}||_{\text{shadow}}^{2}/\epsilon^{2}$, where $\mathcal{O}_{j}$ is the observable to be measured, and $\epsilon$ denotes the desired accuracy.
@@
In 2021, the authors of [119] conducted an experimental implementation of classical shadow, demonstrating its superiority over maximum likelihood in the low sampling regime. This framework’s applicability extends to the prediction of observable polynomials, as discussed in [41]. The strategy involves expressing polynomials as linear combinations of tensor products and substituting the tensor products with independent shadows, enhancing efficiency in probing entanglements in many-body systems. Notably, this method exhibits robustness in the presence of noise, leveraging random unitaries as detailed in [30]. The authors of [75] propose an error-mitigated classical shadow estimation scheme. This result complements [30]. [135] introduced an error-mitigated fermionic classical shadow scheme. Additionally, [65] explores techniques for derandomizing the procedure. In classical computer science, there exists a standard approach to transform a probabilistic argument into a deterministic strategy, maintaining comparable effectiveness without reliance on randomness.
4.2.1 Machine learning with classical shadow